OpenStudy (anonymous):

Name the binomial you can multiply by (x + 9) to get the product .x2+5x-36 a

OpenStudy (pfenn1):

\[(x+9)(x+a)=x^2+5x-36\]What would "a" have to be to ensure \[a+9 = 5\] and \[9a=-36\]?

OpenStudy (anonymous):

awas typo sorry

OpenStudy (lgbasallote):

have you tried \[\large x^2 +5x - 36 \div x+9\]

OpenStudy (lgbasallote):

i mean \[\huge (x^2 + 5x - 36) \div (x+9)\]

OpenStudy (anonymous):

yes i just cant get it rioght

OpenStudy (anonymous):

Can you factor x^2 + 5x - 36

OpenStudy (lgbasallote):

\[\huge \frac{x^2 +5x - 36}{x+9} \leftarrow \text{factor out numerator}\]

OpenStudy (anonymous):

i have a.4-x b.x+6 c.x-4 d.x-6

OpenStudy (anonymous):

and none of them seem to be it

OpenStudy (lgbasallote):

oh it's one of them ;D factor out x^2 + 5x - 36 like @Chlorophyll said

OpenStudy (anonymous):

b

OpenStudy (anonymous):

@joel123 why b? Could you show your calculation here?